# Generalized Spherical Neural Operators: Green’s Function Formulation

### v1

@@ Abstract
Neural operators offer powerful approaches for solving parametric partial differential equations, but extending them to spherical domains remains challenging due to the need to preserve intrinsic geometry while avoiding distortions that break rotational consistency. Existing spherical operators rely on rotational equivariance but often lack the flexibility for real-world complexity. We propose a general operator-design framework based on designable spherical Green’s function and its harmonic expansion, establishing a solid operator-theoretic foundation for spherical learning. Based on this, we propose an absolute and relative position-dependent Green’s function that enables flexible balance of equivariance and invariance for real-world modeling. The resulting operator, Green’s-function Spherical Neural Operator (GSNO) with a novel spectral learning method, can adapt to anisotropic, constraint-rich systems while retaining spectral efficiency. To exploit GSNO, we develop GSHNet, a hierarchical architecture that combines multi-scale spectral modeling with spherical up–down sampling, enhancing global feature representation. Evaluations on diffusion MRI, shallow water dynamics, and global weather forecasting, GSNO and GSHNet consistently outperform state-of-the-art methods. Our results position GSNO as a principled and general framework for spherical operator learning, bridging rigorous theory with real-world complexity.
@@ 1 Introduction
Solving parameterized partial differential equations (PDEs) is a fundamental task across science and engineering. Applications such as weather forecasting, fluid dynamics, and neuroimaging often involve high-order PDEs whose numerical solutions are computationally expensive and intractable. Emerging neural operators provide a promising alternative by approximating solution operators directly from data  (kovachki2024operator). Earlier approaches  (lu2019deeponet; lu2021learning; bhattacharya2021model) learned mapping between function spaces using neural networks but struggled to scale to high-dimensional PDEs. To address this, the Fourier Neural Operator (FNO) (li2020fourier) leverages Fast Fourier transforms (FFTs) to learn in the frequency domain, capturing global patterns and high-frequency modes (kovachki2024operator). This inspired a new generation of operator learning methods and advanced large-scale applications such as high-resolution climate prediction  (pathak2022fourcastnet; liu2024evaluation).
However, FNOs rely on the standard Fourier transform and assume Euclidean geometry. On non-Euclidean manifolds such as the sphere (bonev2023spherical), FFT-based representations introduce distortions: small polar displacements can map to large Cartesian displacements, breaking spatial coherence and degrading performance. To address this, Spherical Fourier Neural Operator (SFNO) is proposed (bonev2023spherical), replacing the FFT with the Spherical Harmonic Transform (SHT). By projecting functions onto spherical harmonic bases, SFNO preserves rotational equivariance on the sphere, ensuring stability under arbitrary input rotations. SFNO-based methods have achieved strong performance on some spherical tasks, e.g., weather prediction  (lin2023spherical; mahesh2024huge1; mahesh2024huge2; hu2025spherical).
@@
(1) Theory: most spherical neural operators are rigorously constructed by Spherical Harmonic Transform and spherical convolution theorem, thereby extending the FNO to the sphere, rather than derived from the integral solution of sphere-native PDEs (bonev2023spherical; lin2023spherical; mahesh2024huge1; mahesh2024huge2; hu2025spherical). This lack of rigorous grounding obscures the physical meaning of the learned integral kernels and limits their generalization and extension in physical systems. (2) Expressiveness: Physical systems involve asymmetric constraints, such as boundary effects, anisotropic forcing, or heterogeneous media (ye2022learning; ye2023locality; lucarini2024detecting; behroozisensitivity), which cannot be captured if operators depend solely on equivariance. However, by underlying design, most spherical operators enforce strict rotational equivariance to gain efficiency for spectral learning, overlooking such asymmetry and thereby limiting their ability to model complex, high-order nonlinear phenomena. (3) Multiscale representations: Most spherical operator networks use spectral parameterizations that capture global structure but are prone to high-frequency truncation and loss of details, making it difficult to preserve fine-grained local patterns  (zhao2019spherical; zhao2021spherical; hu2025spherical).
@@
Figure 1: Method comparison. (Left) Existing approaches use SHT-based spectral learning (Tranditional kernel). (Right) Our method introduces sysyem constraints by designable Green’s function to derive a system-constraint spectral learning (System kernel), yielding system-aware representations.
@@
Our Approach: (1) We first derive the principled formulation of spherical operator learning that begins from a designable Green’s function. This derivation extends the Green’s function naturally from the sphere to the spherical harmonic domain, providing a general operator-theoretic framework of spherical learning, which supports the simulation of different complex systems by designing different Green’s functions. (2) Based on this foundation, we enhance the flexibility of spherical operators on real-world systems by designing an absolute and relative position-dependent Green’s function, deriving the corresponding response to explicitly balance equivariance and invariance. The yielded Green’s-function Spherical Neural Operator (GSNO) balances the efficiency of equivariance with the capacity to capture complex system constraints such as boundaries and distortions. This constraint serves as an explicit inductive bias, enabling GSNO to model heterogeneity in real-world, asymmetric physical systems. (3) Finally, we propose the Generalized Spherical Harmonic Network (GSHNet), a hierarchical architecture that integrates GSNO with multi-scale spectral modeling through spherical up–down sampling, enhancing the expressive power while preserving geometric consistency across resolutions.
Contributions: This work makes the following contributions:
@@
    A general operator-design framework based on the designable spherical Green’s functions (Sec. 4.1).
@@
    We design an absolute and relative position-dependent Green’s function and derive the corresponding response to propose a novel operator (GSNO) that flexibly balance equivariance and invariance, and models complex systems (Sec. 4.2).
@@
    We design GSHNet, a multi-scale spherical network based on GSNO (Sec. 4.3).
@@
    GSNO and GSHNet are evaluated on the diffusion MRI modeling of brain microstructure, spherical shallow water equations and weather prediction, demonstrating consistent improvements in performance over other state-of-the-art models.
@@ 2 Related Work
Geometric equivariance has been explored through transformations such as translation, scaling, and dilation of filters, in both discrete and continuous domains (xu2014scale; sosnovik2021disco; rahman2023truly; chen2023laplace). These approaches contribute to the framework of geometric deep learning by incorporating rich symmetry groups (cohen2016group; bronstein2021geometric). Several methods embed inductive biases to enhance generalization (wad2022equivariance; han2022geometrically); for instance, CNNs are inherently equivariant to translations (li2021survey). Group-equivariant neural networks (cohen2016group; cohen2018spherical), along with spectral approaches such as DISCO (ocampo2022scalable), further exploit symmetries to improve learning efficacy. However, for real-world modeling, there are some components that relax rotational equivariance in most spherical CNN-based models, such as residual pathways, position embeddings, activation functions and local operations (cohen2016group; cohen2018spherical; bonev2023spherical; finzi2021residual; liu2024neural). Other methods are proposed to further introduce complex constraints and relax strict equivariance for better real-world modeling capabilities (finzi2021residual; huang2022equivariant; wang2022approximately; duval2023faenet; pertigkiozoglou2024improving; zheng2024relaxing; li2024physics).
@@
Recent extensions introduce multi-scale learning to enhance feature representation. Some neural operators (li2020multipole; lutjens2022multiscale; raonic2023convolutional; you2024mscalefno; liang2024m; liu2025enhancing) tackle this by combining both upsampling and downsampling operations capturing multiscale information. However, the operations may introduce aliasing artifacts (ronneberger2015u; karras2021alias) and distortions, limiting model accuracy, especially in spherical space (zhao2019spherical; zhao2021spherical).
@@ 3.1 Spherical Harmonic
The spherical harmonic function (muller2006spherical) $Y_{l}^{m}(\theta,\phi)$, with integer degrees $l\geq 0$ and orders $|m|\leq l$, forms an orthonormal basis for square-integrable functions on the unit sphere $S^{2}$. Spherical harmonic transform (SHT) decomposes a function $f\in L^{2}(S^{2})$ into its harmonic coefficients:
@@
Given two functions $f$ and $h$ defined on the sphere $S^{2}$, their spherical convolution is defined as (driscoll1994computing):
@@ 3.2 Spherical Convolution Theorem
To derive the SHT coefficients of a spherical convolution, the result of the spherical convolution theorem (driscoll1994computing) is as follows:
@@
This result demonstrates that spherical convolution in the harmonic domain corresponds to a product of the SHT coefficients of $f$ and $h$ (with $h$ restricted to the $m=0$ mode) in the frequency domain. The full derivation is in the Appendix A. By replacing the classical convolution theorem with this spherical convolution theorem, SFNOs achieve frequency-domain parameterization on the sphere  (bonev2023spherical; lin2023spherical; hu2025spherical). To enhance interpretability, we introduce two complementary SHT-based neural operator derivations in spherical space (Sec. 4.1).
@@ 4.1 Operator Framework via Spherical Green’s Function
The Green’s function method offers a classical strategy to solve PDEs, where the solution is expressed as a convolution integral with the Green’s function as the kernel (li2020neural). Differently, we define $D$, a linear differential operator on the sphere, and consider the following PDE:
$D(g(u))=f(u),\quad u\in S^{2}$ (6)
@@
In this framework, we propose an operator design method based on the designable Green’s function to simulate different systems. For example, classic spectral operators define the mapping on the ideal sphere and assume strict rotational equivariance: Design $G(u,R)$ as $G(R^{-1}u)$. Under this relative-position dependent Green’s function $G(R^{-1}u)$, the prediction target $g(u)$ is given by:
@@
Applying the spherical convolution theorem (Equation 5), the spherical harmonics transform of $g(u)$ becomes:
@@
where $G_{\theta}(l)$ denote the learnable spectral weights parameterized by the neural operator. This consistency with SFNOs verifies the feasibility of this framework of designable Green’s function.
@@
The above derivation relies on spherical convolution theorem. To extend, we propose to parse neural operators entirely from spherical harmonics and present a derivation based on the harmonic extension of Green’s function. Under the square integrability assumption (groemer1996geometric), Green’s function is extended as:
@@
Therefore, this section presents a novel framework to simulate different systems based on designable Green’s functions, deriving the corresponding system kernels, called Green’s Function Formulation, shown in the Figure 1. Equation (11) and (14) are both the solution derived from the relative-position dependent Green’s function, demonstrating two designable or scalable forms of Green’s function: $G(R^{-1}u)$ and its spherical harmonic extension.
Figure 2: The proposed GSNO block (left) and the architecture of GSHNet (right). SHT and ISHT represent spherical harmonic transformation and inverse transformation. Multi-layer perceptrons (MLPs) and two convolutional layers are used for channel interaction.
@@ 4.2 Green’s-function Spherical Neural Operator
Spherical spectral convolutions assume the Green’s function depends solely on the relative position $R^{-1}u$ through rotational equivariance ($G(R,u)=G(R^{-1}u)$). However, this strong assumption ignores the applicability to real-world position-dependent geophysical settings characterized by anisotropy (van2013wu), local heterogeneity (e.g., seismic faults or mantle plumes), and non-periodic boundary conditions (lucarini2024detecting). To address this limitation, the Green’s function is designed by combining the relative and absolute position-dependent terms. This enables modeling of invariant physical properties without sacrificing the efficiency of spectral-domain computation. We define the extended Green’s function hypothesis as:
@@
where $C_{f}$ is the spherical integral of the input function ($f(u)$). The original term $\text{I}({\mathcal{T}_{\text{orig}}})$ and the correction term $\text{I}({\mathcal{T}_{\text{corr}}})$ share $G_{\theta_{1}}^{1}(l)$, and to reduce parameter redundancy between $G_{\theta_{1}}^{1}(l,m)$ and $G_{\theta_{2}}^{2}(l,m)$ and high computational efficiency, we simplify Equation equation 4.2 to:
@@
GSNO combines equivariant symmetry ($G_{1}$) and invariant responses ($G_{2}$), enabling simultaneous modeling of real non-equivariant systems. Specifically, $G_{1}$ encodes more dynamic features, $G_{2}$ explicitly encodes systematic and more stable constraints, including local non-uniformity and boundary constraints (e.g.,anisotropic elasticity tensors, molecular diffusion, weather prediction affected by terrain) while avoiding influence on the feature representation relying on rotation equivariance by sufficient parameter learning. In addition, GSNO  maintains high computational efficiency and low parameters as SFNO. In summary, GSNO generalizes spherical operators by relaxing the strict $SO(3)$ equivariance flexibly, allowing real-world modeling without disrupting spherical geometry.
4.3 Generalized Spherical Harmonic Neural Operator Network
While spectral convolution excels in capturing global dependencies, it is limited in multi-scale modeling. Building on GSNO, we propose a multi-scale spherical harmonic network, GSHNet (Figure 2 right). GSHNet adopts a U-Net structure (zhao2019spherical; zhao2021spherical; hu2025spherical) for expansion and compression of spatial and channel, and incorporates position encoding to enhance global modeling. As a core component, the GSNO block performs scale transformation of the spatial and spectral domains based on SHT and ISHT, and channel transformation via MLP. Specifically, we achieve scale transformation by modifying the number of sampling points along $\theta$ (latitude) and $\phi$ (longitude) in Equation 2 and the degree $l$ in Equation 1, which avoids distortions caused by traditional up- and down-sampling. The downsampling blocks reduce the sampling points and enhance feature expression through MLP, thereby realizing the abstraction of large-scale features. The upsampling restores higher-frequency content, with skip connections providing direct access to high-resolution information. This desgin allows GSHNet  to capture multi-scale interaction via geometric up- and down-sampling on the sphere (details in Appendix B.2).
Figure 3: Predicted geopotential height (H) at 5h and 10h from different methods. Zoom in and provide the residual (Res) to the ground truth for clarity (darker is better).

### v2

@@ Abstract
Neural operators offer powerful approaches for solving parametric partial differential equations, but extending them to spherical domains remains challenging due to the need to preserve intrinsic geometry while avoiding distortions that break rotational consistency. Existing spherical operators rely on rotational equivariance but often lack the flexibility for real-world complexity. We propose a generalized operator-design framework based on designable spherical Green’s function and its harmonic expansion, establishing a solid operator-theoretic foundation for spherical learning. Based on this, we propose an absolute and relative position-dependent Green’s function that enables flexible balance of equivariance and invariance for real-world modeling. The resulting operator, Green’s-function Spherical Neural Operator (GSNO) with a novel spectral learning method, can adapt to non-equivariant systems while retaining spectral efficiency and grid invariance. To exploit GSNO, we develop SHNet, a hierarchical architecture that combines multi-scale spectral modeling with spherical up–down sampling, enhancing global feature representation. Evaluations on diffusion MRI, shallow water dynamics, and global weather forecasting, GSNO and SHNet consistently outperform state-of-the-art methods. The theoretical and experimental results position GSNO as a principled and generalized framework for spherical operator design and learning, bridging rigorous theory with real-world complexity.
@@ 1 Introduction
Solving parameterized partial differential equations (PDEs) is a fundamental task across science and engineering. Applications such as weather forecasting, fluid dynamics, and neuroimaging often involve high-order PDEs whose numerical solutions are computationally expensive and intractable. Emerging neural operators provide a promising alternative by approximating solution operators directly from data  (Kovachki et al., 2024). Earlier approaches  (Lu et al., 2019; 2021; Bhattacharya et al., 2021) learned mapping between function spaces using neural networks but struggled to scale to high-dimensional PDEs. To address this, the Fourier Neural Operator (FNO) (Li et al., 2020a) leverages Fast Fourier transforms (FFTs) to learn in the frequency domain, capturing global patterns and high-frequency modes (Kovachki et al., 2024). This inspired a new generation of operator learning methods and advanced large-scale applications such as high-resolution climate prediction  (Pathak et al., 2022; Liu et al., 2024).
However, FNOs rely on the standard Fourier transform and assume Euclidean geometry. On non-Euclidean manifolds such as the sphere (Bonev et al., 2023), FFT-based representations introduce distortions: small polar displacements can map to large Cartesian displacements, breaking spatial coherence and degrading performance. To address this, Spherical Fourier Neural Operator (SFNO) is proposed (Bonev et al., 2023), replacing the FFT with the Spherical Harmonic Transform (SHT). By projecting functions onto spherical harmonic bases, SFNO preserves rotational equivariance on the sphere, ensuring stability under arbitrary input rotations. SFNO-based methods have achieved strong performance on some spherical tasks, e.g., weather prediction  (Lin et al., 2023; Mahesh et al., 2024a; b; Hu et al., 2025).
@@
(1) Theory: most spherical neural operators are rigorously constructed by Spherical Harmonic Transform and spherical convolution theorem, thereby extending the FNO to the sphere, rather than derived from the integral solution of sphere-native PDEs (Bonev et al., 2023; Lin et al., 2023; Mahesh et al., 2024a; b; Hu et al., 2025). This lack of rigorous grounding obscures the physical meaning of the learned integral kernels and limits their generalization and extension in physical systems. (2) Neural Operator Block: Physical systems involve non-equivariance and asymmetric constraints, such as boundary effects, position-relevant patterns, or heterogeneous media (Ye et al., 2022; 2023; Lucarini and Chekroun, 2024; Behroozi et al., 2025), which cannot be captured if operators depend solely on equivariance. However, by underlying design, most spherical operators enforce strict rotational equivariance to gain efficiency for spectral learning, overlooking such asymmetry and thereby limiting their ability to model complex, high-order nonlinear phenomena. (3) Neural Operator Network: Existing spherical operator networks use ResNet-style architecture Bonev et al. (2023) operating at a single scale to capture global structure, but may limit the expressiveness in modeling climate dynamics (Zhao et al., 2019; 2021; Hu et al., 2025).
@@
Figure 1: Method comparison under the Green’s function framework (Details in Sec. 4.1). (Left) Existing approaches use SHT-based spectral learning (symmetric kernel). (Right) Our method extends sysyem characteristics by designable Green’s function to derive the system-constraint solution.
@@
Our Approach: (1) We first derive the principled formulation of spherical operator learning that begins from a designable Green’s function. This derivation extends the Green’s function naturally from the sphere to the spherical harmonic domain, providing a generalized operator-theoretic framework of spherical learning, which supports the simulation of different complex systems by designing different Green’s functions. (2) Based on this foundation, we enhance the flexibility of spherical operators on real-world systems by designing an absolute and relative position-dependent Green’s function as an explicit inductive bias, deriving the corresponding response to explicitly balance equivariance and invariance. The yielded Green’s-Function Spherical Neural Operator (GSNO) balances the spectral efficiency with the capacity to capture complex system constraints g boundaries and distortions while retaining spherical geometry and grid invariance, enabling the nonlinear modeling in real-world, non-equivariant physical systems. (3) Finally, we propose the Spherical Harmonic Neural Operator Network (SHNet), a hierarchical architecture that integrates GSNO with multi-scale spectral modeling through spherical up–down sampling, enhancing the expressive power while preserving geometric consistency across resolutions.
Contributions: This work mainly makes the following contributions:
@@
    A generalized theoretical framework to derive system-based operator solutions based on the designable spherical Green’s functions (Sec. 4.1).
@@
    We design an absolute and relative position-dependent Green’s function and derive the corresponding response to propose a novel operator (GSNO) that flexibly balances equivariance and invariance and models complex systems while retaining grid invariance(Sec. 4.2).
@@
    We design SHNet, a multi-scale spherical network based on GSNO (Sec. 4.3).
@@
    GSNO and SHNet are evaluated on the diffusion MRI modeling of brain microstructure, spherical shallow water equations and weather prediction, demonstrating consistent improvements in performance over other state-of-the-art models.
@@ 2 Related Work
Geometric equivariance has been explored through transformations such as translation, scaling, and dilation of filters, in both discrete and continuous domains (Xu et al., 2014; Sosnovik et al., 2021; Rahman and Yeh, 2023; Chen et al., 2023). These approaches contribute to the framework of geometric deep learning by incorporating rich symmetry groups (Cohen and Welling, 2016; Bronstein et al., 2021). Several methods embed inductive biases to enhance generalization (Wad et al., 2022; Han et al., 2022); for instance, CNNs are inherently equivariant to translations (Li et al., 2021). Group-equivariant neural networks (Cohen and Welling, 2016; Cohen et al., 2018), along with spectral approaches such as DISCO (Ocampo et al., 2022), further exploit symmetries to improve learning efficacy. However, for real-world modeling, there are some components that relax rotational equivariance in most spherical CNN-based models, e.g., residual pathways, position embeddings, activation functions and local operations (Cohen and Welling, 2016; Cohen et al., 2018; Bonev et al., 2023; Finzi et al., 2021; Liu-Schiaffini et al., 2024). Other methods are proposed to further introduce complex constraints and relax strict equivariance for better real-world modeling capabilities (Finzi et al., 2021; Huang et al., 2022; Wang et al., 2022; Duval et al., 2023; Pertigkiozoglou et al., 2024; Zheng et al., 2024; Li et al., 2024).
@@
Recent extensions introduce multi-scale learning to enhance feature representation. Some neural operators (Li et al., 2020c; Lütjens et al., 2022; Raonic et al., 2023; You et al., 2024; Liang et al., 2024; Liu et al., 2025) tackle this by combining both upsampling and downsampling operations capturing multiscale information. However, the operations may introduce aliasing artifacts (Ronneberger et al., 2015; Karras et al., 2021) and distortions, limiting modeling accuracy, especially in spherical space (Zhao et al., 2019; 2021).
@@ 3.1 Spherical Harmonic
The spherical harmonic function (Müller, 2006) $Y_{l}^{m}(\theta,\phi)$, with integer degrees $l\geq 0$ and orders $|m|\leq l$, forms an orthonormal basis for square-integrable functions on the unit sphere $S^{2}$. Spherical harmonic transform (SHT) decomposes a function $f\in L^{2}(S^{2})$ into its harmonic coefficients:
@@
Given two functions $f$ and $h$ defined on the sphere $S^{2}$, their spherical convolution is defined as (Driscoll and Healy, 1994):
@@ 3.2 Spherical Convolution Theorem
To derive the SHT coefficients of a spherical convolution, the result of the spherical convolution theorem (Driscoll and Healy, 1994) is as follows:
@@
This result demonstrates that spherical convolution in the harmonic domain corresponds to a product of the SHT coefficients of $f$ and $h$ (with $h$ restricted to the $m=0$ mode) in the frequency domain. The full derivation is in the Appendix A. By replacing the classical convolution theorem with this spherical convolution theorem, SFNOs achieve frequency-domain parameterization on the sphere  (Bonev et al., 2023; Lin et al., 2023; Hu et al., 2025). To enhance interpretability, we introduce two complementary SHT-based neural operator derivations in spherical space (Sec. 4.1).
@@ 4.1 Operator Framework via Spherical Green’s Function
The Green’s function method offers a classical strategy to solve PDEs, where the solution is expressed as a convolution integral with the Green’s function as the kernel (Li et al., 2020b). Differently, we define $D$, a linear differential operator on the sphere, and consider the following PDE:
$D(g(u))=f(u),\quad u\in S^{2},$ (6)
@@
In this framework, we propose an operator design method based on the designable Green’s function to simulate different systems. For example, classic spectral operators assume the strict rotational equivariant mapping/system: Design $G(u,R)$ as $G(R^{-1}u)$. Under this relative-position dependent Green’s function $G(R^{-1}u)$, the prediction target $g(u)$ is given by:
@@
Applying the spherical convolution theorem (Equation 5), the spherical harmonic transform of $g(u)$ is simplified as:
@@
where $G_{\theta}(l)$ denote the learnable spectral weights parameterized by the neural operator. This consistency with SFNOs verifies the feasibility of this designable Green’s-function framework.
@@
The above derivation relies on spherical convolution theorem. To extend, we propose to parse neural operators entirely from spherical harmonics and present a derivation based on the harmonic extension of Green’s function. Under the square integrability assumption (Groemer, 1996), Green’s function is extended as:
@@
Therefore, this section presents a comprehensive and solid theoretical framework to simulate different systems based on designable Green’s functions, deriving the corresponding system kernels, shown in the Figure 1. Equation (11) and (14) are both the solution derived from the relative-position dependent Green’s function, demonstrating two designable or scalable forms of Green’s function: $G(R^{-1}u)$ and its spherical harmonic extension.
Figure 2: The proposed GSNO block (left) and the architecture of SHNet (right). SHT and ISHT represent spherical harmonic transformation and inverse transformation. Multi-layer perceptrons (MLPs) and two convolutional layers are used for channel interaction.
@@ 4.2 Green’s-function Spherical Neural Operator
Spherical spectral convolutions assume the Green’s function depends solely on the relative position $R^{-1}u$ through rotational equivariance ($G(R,u)=G(R^{-1}u)$). However, this strong assumption ignores the applicability to real-world non-equivariance, e.g., position-dependent geophysical settings characterized by anisotropy, local heterogeneity, and non-periodic boundary conditions (Van Essen et al., 2013; Lucarini and Chekroun, 2024). To address this limitation, the Green’s function is designed by combining the relative and absolute position-dependent terms. This enables modeling of invariant physical properties without sacrificing the efficiency of spectral-domain computation. We define the extended Green’s function hypothesis as:
@@
where $C_{f}$ is the spherical integral of the input function ($f(u)$). The original term $\text{I}({\mathcal{T}_{\text{orig}}})$ and the correction term $\text{I}({\mathcal{T}_{\text{corr}}})$ share $G_{\theta_{1}}^{1}(l)$. To reduce parameter redundancy between $G_{\theta_{1}}^{1}(l,m)$ and $G_{\theta_{2}}^{2}(l,m)$ and retain high computational efficiency, we simplify Equation 4.2 to:
@@
GSNO combines equivariant response ($G_{1}$) and invariant response ($G_{2}$), enabling simultaneous modeling of real non-equivariant systems. Specifically, $G_{1}$ encodes more dynamic features, $G_{2}$ provides mode-level spectral embedding, which explicitly encodes systematic and more stable constraints, including local non-uniformity and boundary constraints (e.g.,weather prediction affected by terrain) while retaining grid invariance and avoiding influence on the feature representation relying on rotational equivariance by sufficient parameter learning. In addition, GSNO maintains high computational efficiency and low parameters as SFNO. In summary, GSNO improves spherical operators by relaxing the strict $SO(3)$ equivariance flexibly, allowing real-world modeling without disrupting spherical geometry.
4.3 Spherical Harmonic Neural Operator Network
While spectral convolution excels in capturing global dependencies, it is limited in multi-scale modeling. Building on GSNO, we propose a multi-scale spherical harmonic network, SHNet (Figure 2 right). SHNet adopts a U-Net structure (Zhao et al., 2019; 2021; Hu et al., 2025) for expansion and compression of spatial and channel, and incorporates position embedding to enhance global modeling Bonev et al. (2023). As a core component, the GSNO block performs scale transformation of the spatial and spectral domains based on SHT and ISHT, and channel transformation via MLP. Specifically, we achieve scale transformation by modifying the number of sampling points along $\theta$ (latitude) and $\phi$ (longitude) in Equation 2 and the degree $l$ in Equation 1, which avoids distortions caused by traditional up- and down-sampling. The downsampling blocks reduce the sampling points and enhance feature expression through MLP, thereby realizing the abstraction of large-scale features. The upsampling restores higher-frequency content, with skip connections providing direct access to high-resolution information. This desgin allows SHNet  to capture multi-scale interaction via geometric up- and down-sampling on the sphere (details in Appendix B.2).